Ticket Number: HEA-NOC/20041017-2 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: GEANT
Ticket Opened: 20041017 10:28 UTC+1 Problem Start: 20041017 00:00 UTC+1
Ticket Closed: 20041019 17:26 UTC+1 Problem End: 20041019 17:20 UTC+1
The backup link to Geant is down
PRoduction traffic is not affected, but resilience is no longer available
20041017 10:26 BB Investigating
20041017 16:02 BB Visited site - no obvious problems. Will replace cable
between Geant's M40 and HEAnet's Cisco.
20041017 17:47 BB Replaced cable between Geant POP router and HEAnet's
access router, but no change was observed. Have passed diagnostics to Lan
Communications for their analysis.
20041018 10:28 BC Arrange for an onsite ESAT engineer to loop the STM-16
port on phobos. The Rx optical power dropped from -8.8dBm to -32dBm. It
therefore looks as though the Tx laser on the STM-16 card is dead. Asked
Stephen Mulhall of LanComms to arrange for a replacement card to be sent
to City West. Let Geant know that we believe the problem to be with our
20041018 11:32 BC LanComms ticket ref is 2004-SC11249.
20041018 16:16 DC Going onsite at 8 tomorrow morning. Some interruption of
service may be needed when the card is replaced.
20041019 11:13 BC Trouble-shooting on site with DonalC revealed that the
Rx of the Geant interface appears to be the source of the problem. The
replacement Cisco card for phobos should not be necessary, as the HEAnet
interface is working to a hardware loopback - there may have been a
problem yesterday either with the cable or the ESAT-X technician when
trying to test to a hardware loopback.
Geant NOC have accepted that their Rx appears to be at fault and will
arrange for the M40 card to be replaced. Notified ESAT-X via email of the
delivery of the Geant replacement card later today and awaiting the name
of the Dante/Geant contact who will install the card.
20041019 17:23 BC Geant's engineer arrived on site and replaced the line
card in the M40. The line came back up and is running clean. Will leave
ticket open until tomorrow.
This ticket can be monitored at http://www.hea.net/tickets/20041017-2